OpenStudy (princeevee):

solve for w p=2L+2w w=p-2L/2 w=p-L w=2L-p/2 w=p-2L/L

imqwerty (imqwerty):

A

OpenStudy (anonymous):

\[p=2L+2w\] \[p-2L=2w\] \[\frac{ p-2L }{ 2 }=w\]